The vapour density of a mixture A ( molecular mass=40)and gas B (mm=80) is 25.find mole% of gas B in the mixture

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
 
Avg.. molar mass = 25/2
M = (n1*40 + n2*80)/(n1 + n2) = 25/2 
=> 80n1 + 160n2 = 25n1 + 25n2
=> 55n1 + 135n2 = 0
=> 11n1 + 27n2 = 0
n1/n2 = + 27/11..............(1)
as mol can’t be negative.
mol % of n2 = n2/(n1+n2) * 100
                   = 1/(n1/n2  + 1)*100
                   =1/(1 + 27/11) * 100
                   = 11/38 * 100
                   = 28.9%
 
Hope it helps
 
Regards
 
Arun (askIITians forum expert)